OpenStudy (anonymous):

solve integrate x^(-1/2)sin x dx

OpenStudy (anonymous):

\[\int\limits_{}^{}x ^{^{-1/2}}(\sin x )dx\]

OpenStudy (lgbasallote):

are you allowed to use integration by parts?

OpenStudy (anonymous):

ya allow by using any method..

OpenStudy (lgbasallote):

i think IPB is the only way though =)) are you familiar with LIATE? (a method on which you'll use as u and which as dv)

OpenStudy (lalaly):

this is not an elementary fnction,, so by parts wont work

OpenStudy (lgbasallote):

aww that's a downer :(

OpenStudy (anonymous):

i hv tried the intergrate by part but the power keep increase can't eliminate the intergral

OpenStudy (lalaly):

how far in integration methods have u got?

OpenStudy (anonymous):

i think three 1st by formular 2nd by parts 3rd by subtitution

OpenStudy (lgbasallote):

what about algebraic sub?

OpenStudy (lalaly):

are u sure u wrote the question right?

OpenStudy (lgbasallote):

wont algebraic sub work? let u = \(\sqrt x\) \(u^2 = x\) 2udu = dx \(\LARGE \int \frac{\sin (u^2) 2\cancel{u}du}{\cancel{u}}\) \(\LARGE 2\int \sin (u^2) du\)

OpenStudy (anonymous):

i just copy from my friend i dont know whether it is correct or nt

OpenStudy (anonymous):

thn can you solve the sin(x^2)??

OpenStudy (lalaly):

ok lg how can u solve sin(u^2)?

OpenStudy (lgbasallote):

i was hoping you can do it ~''~

OpenStudy (lalaly):

it cant be done using the methods hes taken

OpenStudy (lalaly):

recheck the question

OpenStudy (lgbasallote):

i wolframed it and a fresnel thingy came out...i think this is a syntax problem

OpenStudy (anonymous):

i think cant do this question.. i try to recheck the question with my friend.. anyway thx for helping me.. =)